# A Comparative Study on State-Action Spaces for Learning Viewpoint Selection and Manipulation with Diffusion Policy

### v2

@@ Abstract
Robotic manipulation tasks often rely on static cameras for perception, which can limit flexibility, particularly in scenarios like robotic surgery and cluttered environments where mounting static cameras is impractical. Ideally, robots could jointly learn a policy for dynamic viewpoint and manipulation. However, it remains unclear which state-action space is most suitable for this complex learning process. To enable manipulation with dynamic viewpoints and to better understand impacts from different state-action spaces on this policy learning process, we conduct a comparative study on the state-action spaces for policy learning and their impacts on the performance of visuomotor policies that integrate viewpoint selection with manipulation. Specifically, we examine the configuration space of the robotic system, the end-effector space with a dual-arm Inverse Kinematics (IK) solver, and the reduced end-effector space with a look-at IK solver to optimize rotation for viewpoint selection. We also assess variants with different rotation representations. Our results demonstrate that state-action spaces utilizing Euler angles with the look-at IK achieve superior task success rates compared to other spaces. Further analysis suggests that these performance differences are driven by inherent variations in the high-frequency components across different state-action spaces and rotation representations.
@@ I Introduction
Figure 1: Visuomotor policies control manipulation and viewpoint selection with different state-action spaces. The Look-At IK automatically determines the orientation of the viewpoint arm end-effector. Green and blue arrows represent the local $y$ and $z$ axes $\mathbf{w}_{y}$ and $\mathbf{w}_{z}$ of the viewpoint end-effector, respectively. The red dot indicates a distant point $\mathbf{p_{f}}$ along the orange line, showing the direction of $\mathbf{w}_{z}$. The pink star marks the visual target $\mathbf{p_{t}}$, set with a fixed offset from $\mathbf{p_{m}}$.
@@
These challenges highlight the need for simultaneous viewpoint selection and manipulation within a robotic system featuring a dynamic camera. Ideally, robots could jointly learn viewpoint selection and manipulation, allowing them to generalize patterns of what, when, and how to move, along with determining when and from where to update their understanding of the environment to complete tasks. However, integrating manipulation with viewpoint selection introduces complexity, and there is no consensus on the choice of state-action space for this joint learning process. Therefore, in this work, we explore how different state-action spaces impact the performance of visuomotor policies that combine viewpoint selection with manipulation. As illustrated in Fig. 1, we conduct a comparative study by evaluating diffusion policies learned from different state-action spaces, including the configuration space of the robotic system, the end-effector space with a dual-arm IK solver §IV-A, and the reduced end-effector space that we refer to as the look-at space §IV-B. For the look-at space, we design a look-at IK solver that optimizes the camera rotation based on a viewpoint selection objective. This reduces the dimensionality of the state-action space, which enhances the efficiency of diffusion policy training and provide a more effective comparison with the standard end-effector space. We also examine variants of the end-effector space and the look-at space using different rotation representations.
For this comparative study, we developed a simulation environment for the dual-arm robotic system with a dynamic viewpoint and performed various experiments in simulation and real-world environments. Our results suggest that the learned diffusion policy, integrated with the look-at IK solver, achieves a higher task success rate (§V-B, SV-C). Although usually high dimensionality is considered as the primary factor for affecting policy learning [10, 21, 1], further analysis indicates that performance differences among diffusion policies with various state-action spaces might be more likely attributed to differences in High Frequency Component (HFC) in these spaces with different rotation representation (§V-D) instead of the dimensionality. We discuss the implications of this analysis concerning the spatial and optimization properties of different rotation representations and findings from the machine learning community on learning from HFC in commonly used diffusion model backbones (§VI-A). Finally, we conclude with future work for searching diffusion backbones that better capture HFC (§VI-B). Our contributions are summarized as follows:
  • A comparative study of diffusion policies for viewpoint selection and manipulation learned from different state-action spaces for various tasks.
  • Evidence of the correlation between the HFC on different state-action spaces and the performance of the learned visuomotor policies.
@@
  • Open-source implementations of our proposed method and the simulation environment to facilitate future research on viewpoint selection and manipulation. [^1] https://github.com/Apollo-Lab-Yale/spaces_comparative_study
@@ II Related Works
II-A Viewpoint Selection and Manipulation
@@
Research has shown that viewpoint selection can improve the success rate of manipulation tasks in cluttered and occluded environments [9, 16]. Most optimization-based planners with dynamic viewpoint select the next best view based on metrics of information gain, such as the number of revealed occluded voxels [4], or the reduction in grasp quality uncertainty [24]. Recent work has started applying learning to viewpoint selection and manipulation. For instance, Saito et al. [31] used a recurrent neural network for a 6-DOF manipulation scenario, though they decoupled viewpoint selection and manipulation into separate phases. Lv et al. [22] utilized model-based reinforcement learning to train a policy for dual 7-DOF arms, showing that learning-based viewpoint selection can enhance task execution compared to static viewpoints. However, this work is limited to short-horizon tasks without any insight on impact of different state-action spaces under such complex decision making process.
II-B Robot Learning and Diffusion Models
With the rise of large generative models, recent advances in IL have integrated more sophisticated network architectures, including transformer-based models [33, 44] and diffusion-based policies [5, 46]. Diffusion policies, inspired by generative models for images and videos that employ a denoising process [14, 27], typically utilize Convolutional Neural Networks (CNNs), such as UNet [30], or Transformers [41] as their backbone. These models have shown significant potential and have been successfully applied to various robotic manipulation tasks [32, 20]. However, diffusion policies have not yet been explored in the context of manipulation with viewpoint selection, and while some studies have examined the connection between state-action spaces and data efficiency [48, 8], there is limited analysis on how different state-action spaces influence performance from a frequency perspective. In contrast, the machine learning community has studied how common diffusion backbones interact with frequency components in datasets. For instance, residual structures in CNNs, often used in UNet backbones, can suppress learning from HFC [43], whereas Vision Transformers are more effective at capturing low-frequency components (LFC) and are less suited for HFC [25]. Building on these insights, we aim to analyze the performance of different state-action spaces in robotic tasks through the lens of frequency analysis.
@@ III-A Problem Formulation
This work employs imitation learning to train a visuomotor policy for a 17-DOF dual-arm robotic system shown in Fig. 2. This process requires a demonstration dataset $\mathcal{D}=\{r_{n}\}_{n=1}^{N}$, which contains $N$ demonstration rollouts $r_{n}$. Each rollout is a sequence of observation-action pairs $(\mathbf{O}_{t},\mathbf{A}_{t})$ at each timestep $t$. This dataset is collected by a human expert with an intrinsic policy $\pi^{*}$. The goal of imitation learning is to mimic $\pi^{*}$ by learning a policy $\pi$ using $\mathcal{D}$ that maps observations $\mathbf{O}_{t}$ to actions $\mathbf{A}_{t}$:
@@
In our study, $\mathbf{O}_{t}$ consists of $\mathbf{v}_{t}$, the RGB image from a camera on the viewpoint arm, and $\mathbf{s}_{t}$, the state of the robot in a space for policy learning.
@@ III-B State-Action Spaces
We explore different state-action spaces, including configuration space, end-effector space, and look-at space.
Configuration Space (C-Space) is a multidimensional space where each dimension corresponds to a DOF of the robot. The C-Space $\mathcal{C}$ encompasses all possible configurations of a robotic system. For our dual-arm system:
@@
End-Effector Space refers to the set of all positions and orientations that end-effectors can reach within the workspace. For our dual-arm system, the end-effector space $\mathcal{E}=\mathcal{E}_{m}\times\mathcal{E}_{v}$ combines the manipulation arm end-effector space $\mathcal{E}_{m}$ and the viewpoint arm end-effector space $\mathcal{E}_{v}$. The end-effector vector of single arm consists of a position vector $\mathbf{p}=[x,y,z]^{T}\in\mathbb{R}^{3}$ and a rotation representation. We consider three rotation representations: Euler angles $\mathbf{R}_{\text{E}}\in\mathbb{R}^{3}$, unit quaternions $\mathbf{R}_{\text{Q}}\in\mathbb{H}_{1}$, and axis-angle $\mathbf{R}_{\text{AA}}\in(\mathbb{R}^{3},\mathbb{R})$. For Euler angles, $\mathbf{R}_{\text{E}}=[\alpha,\beta,\gamma]^{T}$, where $\alpha$, $\beta$, and $\gamma$ denote roll, pitch, and yaw angles. For unit quaternions, $\mathbf{R}_{\text{Q}}=[q_{0},q_{1},q_{2},q_{3}]^{T}$, with $q_{0}$ as the scalar part, $[q_{1},q_{2},q_{3}]^{T}$ as the vector part, and $q_{0}^{2}+q_{1}^{2}+q_{2}^{2}+q_{3}^{2}=1$. For axis-angle, $\mathbf{R}_{\text{AA}}=(\mathbf{u},\theta)$, where $\mathbf{u}=[u_{x},u_{y},u_{z}]^{T}$ is a unit vector representing the rotation axis, and $\theta$ is the rotation angle in radian.
When using end-effector space, both arms in the dual-arm system use the same rotation representation. We consider three variants of the dual-arm end-effector space, including End-Effector Euler Space $\mathcal{E}_{\text{E}}=\{[\mathbf{p}_{m},\mathbf{R}_{\text{E},m},\mathbf{p}_{v},\mathbf{R}_{\text{E},v}]\}$, End-Effector Quaternion Space $\mathcal{E}_{\text{Q}}=\{[\mathbf{p}_{m},\mathbf{R}_{\text{Q},m},\mathbf{p}_{v},\mathbf{R}_{\text{Q},v}]\}$, and End-Effector Axis-Angle Space $\mathcal{E}_{\text{AA}}=\{[\mathbf{p}_{m},\mathbf{R}_{\text{AA},m},\mathbf{p}_{v},\mathbf{R}_{\text{AA},v}]\}$. $\mathbf{p}_{m}$ and $\mathbf{p}_{v}$ are the respective positions of the manipulation and viewpoint end-effectors, $\mathbf{R}_{\text{E},m}$ and $\mathbf{R}_{\text{E},v}$ are the respective Euler angles of the manipulation and viewpoint end-effectors, $\mathbf{R}_{\text{Q},m}$ and $\mathbf{R}_{\text{Q},v}$ are the respective quaternions of the manipulation and viewpoint end-effectors, and $\mathbf{R}_{\text{AA},m}$ and $\mathbf{R}_{\text{AA},v}$ are the respective axis-angle representations of the manipulation and viewpoint end-effectors.
@@
Look-At Space simplifies the end-effector space by excluding the orientation of the viewpoint arm, which is automatically determined (see §IV-B) [29]. Similar to the end-effector space, we consider three variants of the look-at space: Look-At Euler Space $\mathcal{L}_{\text{E}}=\{[\mathbf{p}_{m},\mathbf{R}_{\text{E},m},\mathbf{p}_{v}]\}$, Look-At Quaternion Space $\mathcal{L}_{\text{Q}}=\{[\mathbf{p}_{m},\mathbf{R}_{\text{Q},m},\mathbf{p}_{v}]\}$, and Look-At Axis-Angle Space $\mathcal{L}_{\text{AA}}=\{[\mathbf{p}_{m},\mathbf{R}_{\text{AA},m},\mathbf{p}_{v}]\}$ after excluding $\mathbf{R}_{\text{E},v}$, $\mathbf{R}_{\text{Q},v}$, or $\mathbf{R}_{\text{AA},v}$.
@@ IV Technical Details
IV-A Dual-Arm IK
The dual-arm IK solver is designed to convert from any end-effector space variant $\mathcal{E}_{\text{E}},\mathcal{E}_{\text{Q}},\mathcal{E}_{\text{AA}}$ to the C-space $\mathcal{C}$. The goal is to find the joint configurations $\mathbf{c}$ that minimize the transformation error between the desired and current positions $\mathbf{p}_{\text{m}}^{*},\mathbf{p}_{\text{m}}$ and orientations $\mathbf{R}_{\text{m}}^{*},\mathbf{R}_{\text{m}}$ of the manipulation end-effector, and similarly for the viewpoint end-effector.

### v3

# Optimizing Active Perception for Learning Simultaneous Viewpoint Selection and Manipulation with Diffusion Policy

@@ Abstract
Robotic manipulation tasks often rely on static cameras for perception, which can limit flexibility, particularly in scenarios like robotic surgery and cluttered environments where mounting static cameras is impractical. Ideally, robots could jointly learn a policy for dynamic viewpoint and manipulation. However, dynamic viewpoint control requires additional degrees of freedom and intricate coordination with manipulation, which results in more challenging policy learning than single-arm manipulation. To address this complexity, we propose an integrated learning framework that combines diffusion policy with a novel look-at inverse kinematics solver for active perception. Our framework helps better coordinating between perception and manipulation. It automatically optimizes camera orientation for viewpoint selection, while allowing the policy to focus on essential manipulation and positioning decisions. We demonstrate that our integrated approach achieves superior performance and learning efficiency compared to directly applying diffusion policies to configuration space or end-effector space with various rotation representations. Further analysis suggests that these performance differences are driven by inherent variations in the high-frequency components across different state-action spaces.
@@ I Introduction
Figure 1: An overview of the visuomotor policy learning pipeline and a comparison of the state-action spaces. While policies can be trained in Configuration Space (C-Space) or with a standard Dual-Arm IK, our proposed Look-At IK automatically determines the orientation of the viewpoint arm end-effector. Green and blue arrows represent the local $y$ and $z$ axes $\mathbf{w}_{y}$ and $\mathbf{w}_{z}$ of the viewpoint end-effector, respectively. The red dot indicates a distant point $\mathbf{p_{f}}$ along the orange line, showing the direction of $\mathbf{w}_{z}$. The pink star marks the visual target $\mathbf{p_{t}}$, set with a fixed offset from $\mathbf{p_{m}}$.
@@
These challenges highlight the need for simultaneous viewpoint selection and manipulation within a robotic system featuring a dynamic camera. Ideally, robots could jointly learn viewpoint selection and manipulation, allowing them to generalize patterns of what, when, and how to move, along with determining when and from where to update their understanding of the environment to complete tasks. However, integrating manipulation with viewpoint selection introduces significant difficulty for visuomotor policy learning due to the increased complexity and intricate coordination required between the camera and the gripper.
To address this challenge, we introduce a novel framework that synergistically combines a visuomotor diffusion policy with a specialized look-at Inverse Kinematics (IK) solver, as illustrated in Fig. 1. Our look-at IK solver, detailed in §IV-B, decouples the learning problem by programmatically optimizing the camera’s orientation to maintain focus on the manipulation target. This approach defines the look-at space, which is a new, reduced end-effector space that significantly simplifies the learning task for the policy. By handling the complex camera orientation automatically, our framework allows the diffusion policy to focus on the essential Degrees of Freedom (DOF) for the task, leading to improved training efficiency and superior performance. To validate the effectiveness of our framework, we demonstrate its performance against policies trained in more conventional state-action spaces, including the configuration space, full end-effector space and its variants with different rotation representations.
To facilitate evaluation, we developed new simulation environments for dual-arm active perception systems. Our experiments in both simulation and the real world demonstrate that our integrated framework achieves a significantly higher task success rate and better training efficiency (§V-B, §V-C). Crucially, our analysis reveals that the performance gains are not solely attributable to the reduced dimensionality, which is a common assumption in policy learning [9, 19, 1]. Instead, we find a strong correlation between policy performance and the prevalence of High-Frequency Components (HFC) within the state-action space (§V-D). Our proposed look-at space exhibits lower HFC, making the learning problem more tractable for the diffusion model architecture. This insight provides a deeper understanding of the factors affecting visuomotor policy performance in complex coordination tasks. In summary, our contributions are:
  • A novel framework that integrates a diffusion policy with a look-at IK solver to automate optimal viewpoint orientation, significantly improving performance and training efficiency for manipulation tasks with active perception.
  • New simulation environments for developing and evaluating dual-arm active perception and manipulation systems.
  • An analysis demonstrating that the performance of visuomotor policies is strongly correlated with the HFC of the state-action space, offering a new perspective beyond just dimensionality.
@@
  • Open-source implementation of our framework to facilitate future research on viewpoint selection and manipulation.  [^1] https://github.com/Apollo-Lab-Yale/spaces_comparative_study
@@ II Related Works
II-A Robot Learning and Diffusion Models
With the rise of large generative models, recent advances in IL have integrated more sophisticated network architectures, including transformer-based models [30, 41] and diffusion-based policies [5, 44]. Diffusion policies, inspired by generative models for images and videos that employ a denoising process [12, 23], typically utilize Convolutional Neural Networks (CNNs), such as UNet [26], or Transformers [38] as their backbone. These models have shown significant potential and have been successfully applied to various robotic manipulation tasks [28, 18].
Concurrent work has also begun to apply diffusion policies to manipulation with active perception. Xiong et al. [43] utilized an end-to-end diffusion policy that directly operates in the full end-effector space to learn active perception and manipulation. In contrast, our framework introduces the look-at IK solver as a structural prior to explicitly decouple camera orientation from the learned policy, which significantly improves both performance and training efficiency.
II-B Viewpoint Selection and Manipulation
@@
Research has shown that viewpoint selection can improve the success rate of manipulation tasks in cluttered and occluded environments [8, 14]. Most optimization-based planners with dynamic viewpoint select the next best view based on metrics of information gain, such as the number of revealed occluded voxels [4], or the reduction in grasp quality uncertainty [21]. Recent work has started applying learning to viewpoint selection and manipulation. For instance, Saito et al. [27] used a recurrent neural network for a 6-DOF manipulation scenario, though they decoupled viewpoint selection and manipulation into separate phases. Lv et al. [20] utilized model-based reinforcement learning to train a policy for dual 7-DOF arms, showing that learning-based viewpoint selection can enhance task execution compared to static viewpoints. However, this work is limited to short-horizon tasks and without any simplification on the state-action space from the perspective of kinematics.
@@ III-A Problem Formulation
Our proposed framework employs IL to train a visuomotor policy for a 17-DOF dual-arm robotic system shown in Fig. 2. This process requires a demonstration dataset $\mathcal{D}=\{r_{n}\}_{n=1}^{N}$, which contains $N$ demonstration rollouts $r_{n}$. Each rollout is a sequence of observation-action pairs $(\mathbf{O}_{t},\mathbf{A}_{t})$ at each timestep $t$. This dataset is collected by a human expert with an intrinsic policy $\pi^{*}$. The goal of imitation learning is to mimic $\pi^{*}$ by learning a policy $\pi$ using $\mathcal{D}$ that maps observations $\mathbf{O}_{t}$ to actions $\mathbf{A}_{t}$:
@@
In our framework, $\mathbf{O}_{t}$ consists of $\mathbf{v}_{t}$, the RGB image from a camera on the viewpoint arm, and $\mathbf{s}_{t}$, the state of the robot in a space for policy learning.
@@ III-B State-Action Spaces
A key design choice of our framework is the look-at space, which simplifies the learning problem. By programmatically optimizing the camera’s orientation, it decouples viewpoint rotation from the policy’s action space, allowing the policy to focus on essential manipulation and positioning tasks.
To rigorously validate this approach, we evaluate our method against two standard baselines: Configuration Space and End-Effector Space. These serve as strong points of comparison due to their use in diffusion policies for manipulation [5, 29], including concurrent work on active perception [43]. For a comprehensive analysis, we also evaluate variants of the end-effector and look-at spaces using different rotation representations.
This subsection details these spaces in order of increasing abstraction, from the robot’s native Configuration Space, to the task-centric End-Effector Space, and concluding with our proposed Look-At Space.
Configuration Space (C-Space) is the foundational representation of a robot’s state, directly characterizing its joint angles. As a multidimensional space where each dimension corresponds to a DOF, $\mathcal{C}$ encompasses all possible configurations of the system. For our dual-arm robot:
@@
End-Effector Space offers a more intuitive, task-oriented representation by defining the positions and orientations of the end-effectors in the workspace. States and actions in this space are mapped back to C-Space using a standard IK solver. For our dual-arm system, the end-effector space $\mathcal{E}=\mathcal{E}_{m}\times\mathcal{E}_{v}$ combines the manipulation arm end-effector space $\mathcal{E}_{m}$ and the viewpoint arm end-effector space $\mathcal{E}_{v}$. The end-effector vector of single arm consists of a position vector $\mathbf{p}=[x,y,z]^{T}\in\mathbb{R}^{3}$ and a rotation representation. We consider three rotation representations: Euler angles $\mathbf{R}_{\text{E}}\in\mathbb{R}^{3}$, unit quaternions $\mathbf{R}_{\text{Q}}\in\mathbb{H}_{1}$, and axis-angle $\mathbf{R}_{\text{AA}}\in(\mathbb{R}^{3},\mathbb{R})$. For Euler angles, $\mathbf{R}_{\text{E}}=[\alpha,\beta,\gamma]^{T}$, where $\alpha$, $\beta$, and $\gamma$ denote roll, pitch, and yaw angles. For unit quaternions, $\mathbf{R}_{\text{Q}}=[q_{0},q_{1},q_{2},q_{3}]^{T}$, with $q_{0}$ as the scalar part, $q_{1},q_{2},q_{3}]^{T}$ as the vector part, and $q_{0}^{2}+q_{1}^{2}+q_{2}^{2}+q_{3}^{2}=1$. For axis-angle, $\mathbf{R}_{\text{AA}}=(\mathbf{u},\theta)$, where $\mathbf{u}=[u_{x},u_{y},u_{z}]^{T}$ is a unit vector representing the rotation axis, and $\theta$ is the rotation angle in radians.
Based on the single-arm end-effector representation, we consider three variants of the dual arm end effector space, including the Euler space of the end effector $\mathcal{E}_{\text{E}}=\{[\mathbf{p}_{m},\mathbf{R}_{\text{E},m},\mathbf{p}_{v},\mathbf{R}_{\text{E},v}]\}$, the quaternion space of the end effector $\mathcal{E}_{\text{Q}}=\{[\mathbf{p}_{m},\mathbf{R}_{\text{Q},m},\mathbf{p}_{v},\mathbf{R}_{\text{Q},v}]\}$, and the axes angle space of the end effector $\mathcal{E}_{\text{AA}}=\{[\mathbf{p}_{m},\mathbf{R}_{\text{AA},m},\mathbf{p}_{v},\mathbf{R}_{\text{AA},v}]\}$. $\mathbf{p}_{m}$ and $\mathbf{p}_{v}$ are the respective positions of the manipulation and viewpoint end-effectors, $\mathbf{R}_{\text{E},m}$ and $\mathbf{R}_{\text{E},v}$ are the respective Euler angles of the manipulation and viewpoint end-effectors, $\mathbf{R}_{\text{Q},m}$ and $\mathbf{R}_{\text{Q},v}$ are the respective quaternions of the manipulation and viewpoint end-effectors, and $\mathbf{R}_{\text{AA},m}$ and $\mathbf{R}_{\text{AA},v}$ are the respective axis-angle representations of the manipulation and viewpoint end-effectors.
@@
Look-At Space is a principled simplification of the end-effector space for active perception and manipulation scenarios by excluding the orientation of the viewpoint arm, which is automatically determined (see §IV-B) [25]. Similar to the end-effector space, we consider three variants of the look-at space: Look-At Euler Space $\mathcal{L}_{\text{E}}=\{[\mathbf{p}_{m},\mathbf{R}_{\text{E},m},\mathbf{p}_{v}]\}$, Look-At Quaternion Space $\mathcal{L}_{\text{Q}}=\{[\mathbf{p}_{m},\mathbf{R}_{\text{Q},m},\mathbf{p}_{v}]\}$, and Look-At Axis-Angle Space $\mathcal{L}_{\text{AA}}=\{[\mathbf{p}_{m},\mathbf{R}_{\text{AA},m},\mathbf{p}_{v}]\}$ after excluding $\mathbf{R}_{\text{E},v}$, $\mathbf{R}_{\text{Q},v}$, or $\mathbf{R}_{\text{AA},v}$.
@@
III-C Diffusion Policy
Our framework uses a diffusion policy [5] as the visuomotor component, which is trained by IL within the state-action spaces described previously. We use U-Net [26] as the backbone for the diffusion policy. The diffusion policy employs receding horizon control and models the conditional distribution $p(\mathbf{A}_{t}\mid\mathbf{O}_{t})$. During inference, the diffusion policy starts from Gaussian noise $\mathbf{A}_{t}^{K}$, where $K$ is the number of steps for denoising. The denoising process repeatedly utilizes the following equation:
until the process reaches the output $\mathbf{A}_{t}^{0}$ of the policy.
@@ IV Technical Details
This section details the IK solvers that map the task-level spaces back to the C-Space. We first describe the standard End-Effector IK, which serves as the baseline and foundation for our method. We then present our proposed Look-At IK solver, which introduces a novel optimization objective specifically designed to automate viewpoint control for active perception with manipulation tasks.
IV-A End-Effector IK
A traditional end-effector IK solver, when applied to our dual-arm system, is designed to convert from any end-effector space variant $\mathcal{E}_{\text{E}},\mathcal{E}_{\text{Q}},\mathcal{E}_{\text{AA}}$ to the C-space $\mathcal{C}$. The goal is to find the joint configurations $\mathbf{c}$ that minimize the transformation error between the desired and current positions $\mathbf{p}_{\text{m}}^{*},\mathbf{p}_{\text{m}}$ and orientations $\mathbf{R}_{\text{m}}^{*},\mathbf{R}_{\text{m}}$ of the manipulation end-effector, and similarly for the viewpoint end-effector.